### v2

# SwapMoE: Efficient Memory-Constrained Serving of Large Sparse MoE Models via Dynamic Expert Pruning and Swapping

Abstract.
Mixture of experts (MoE) is a popular technique to improve capacity of large models with conditionally-activated parallel neural network modules (experts). Due to its remarkable scaling performance with sparse computation, it is widely used in modern Large Language Models (LLMs) and Large Vision Models (LVMs). However, serving such large models on edge devices is challenging due to memory constraints. Typical solutions like memory swapping or weight pruning may lead to significantly higher latency or severe accuracy loss.
In this paper, we introduce SwapMoE, a framework for efficient continuous MoE-based large models serving with tunable memory budgets. The main idea of SwapMoE is to keep a small dynamic set of important experts, namely Virtual Experts, in the main memory for inference, while seamlessly maintaining how the Virtual Experts map to the actual experts. We use a profiling-guided planner to allocate the resources for SwapMoE that can fully utilize the memory budgets and bandwidth, and an importance-aware scheduler to efficiently identify, update, and use the Virtual Experts for accurate inference.
To evaluate SwapMoE, we conduct experiments on multiple edge devices with state-of-the-art MoE-based Large Language Models and Large Vision Models. The results demonstrate remarkable performance of SwapMoE under various memory constraints. Specifically, SwapMoE can enable running large MoE models under tight memory budgets with similar latency to pruned compact models, while with significantly higher accuracy.
submissionid: #399copyright: acmcopyright
1. Introduction
Recently, the world has witnessed the great advancement of deep learning (DL), represented by pre-trained large Transformers. Such an advancement is driven by the phenomenon that the larger model capacity generally leads to higher intelligence (Kaplan et al., 2020). Among various attempts to scale up neural networks, Mixture of Experts (MoE) is a promising technique that can avoid linearly increasing computation based on conditional sparse computation.
An MoE model can be created from a normal model by replacing several layers with MoE layers, each consisting of multiple parallel neural network modules (experts). In each layer, the experts are trained to accommodate different data distributions, and only the most relevant experts are activated (by a gate module) and used for computation for each input sample during the inference process. Due to this benefit, MoE is now widely used in almost all AI applications, including computer vision (Riquelme et al., 2021; Li et al., 2023; Xue et al., 2021), neural language processing (Fedus et al., 2022; Shazeer et al., 2017), and speech recognition (You et al., 2021, 2022).
However, unlike cloud servers that can elastically expand the memory with distributed computing, edge devices often have limited memory, which prevent them to enjoy the benefits of MoE. For example, a Switch Transformer (Fedus et al., 2022) with 64 experts per layer requires 14 GiB of memory for inference, which is impossible to fit in edge devices that typically have only 8 GiB, 4 GiB, or lower memory size. What’s more, many edge devices need to serve multiple applications, and the system-allocated memory budget for each application is even more limited.
Researchers have proposed various methods to reduce the memory consumption of MoE models. The most straightforward approach from the system perspective is to use memory swapping, i.e. dynamically loading/unloading the expert parameters from/to the external memory. For example, Huang et al. (2023) propose to load experts into the memory on demand to reduce memory footprint, EdgeMoE (Yi et al., 2023) focuses on memory swapping in the decoding phase of MoE-based language models and reduces the swapping overhead using quantization. Brainstorm  (Cui et al., 2023) proposed consolidating related experts onto the same device to reduce communication overhead and combining multiple expert kernels into a single kernel for computation, albeit without reducing memory usage. There are also algorithm-perspective approaches (Chen et al., 2022; Kim et al., 2021) that propose to prune the experts to reduce model size. However, swapping-based approaches either have to trade latency for a reduced memory footprint, while pruning-based methods would lead to accuracy loss and require model training. How to efficiently serve an off-the-shelf MoE model under memory constraints remains challenging.
Fortunately, there is a unique property in many MoE workloads that can be exploited for performance optimization - data distribution locality. For example, in generative language models, the output tokens are produced one by one, and they mostly belong to the same semantic context. The successive queries of a user are also semantically related. In object detection, the input images fed into the model are usually continuous video frames. Such locality can potentially lead to consistent activation patterns of MoE experts, therefore create room for more intelligent expert management. Furthermore, the temporal locality property of video analytics has been extensively studied in prior work (Wakatsuki et al., 2021) for accelerating normal CNN models, while how it can be used in MoE for memory reduction is unknown yet.
Our work. We introduce SwapMoE to enable efficient continuous MoE serving under memory constraints. The key idea is to maintain a dynamically-updated compact set of Virtual Experts in the main memory for MoE inference, instead of the redundant set of all experts in the original model. The weights of Virtual Experts are seamlessly updated according to the data distribution locality and profiled hardware capabilities. As such, the memory footprint and latency of each individual inference process in SwapMoE are the same as running a smaller MoE model, while the advantage of large-capability MoE remains since each expert still has the chance to participate in the computation.
Achieving efficient and accurate inference under memory constraint with SwapMoE requires to further address two main challenges. The first is to dynamically identify the important experts as the Virtual Experts to use in the main memory. The second is to seamlessly update the Virtual Experts to avoid extra I/O waiting latency and memory consumption.
We adopt a two-phase design, including an online model serving phase and an offline planning phase, and address the aforementioned challenges with various techniques in both phases. In the online phase, the job of SwapMoE is to handle the whole inference process with the small set of Virtual Experts, update the Virtual Experts to keep the most important experts in memory, and make sure the update process is non-blocking. The main difficulty to maintain the Virtual Experts is to infer the importance of each expert (including the off-memory ones) without using the expert for actual computation. We find a lightweight indicator based on the understanding of MoE operations. During inference, we modify the gating logic to redirect the control flows from the original experts to the Virtual Experts, and update the Virtual Experts with lightweight, amortized, asynchronous parameter loading.
The goal of the offline phase is to plan the execution strategy of SwapMoE (including the per-layer memory allocation and expert update frequency) to meet the memory constraints. The total number of experts in Virtual Experts is determined by the memory budget, but how to divide the budget to different layers is based on each layer’s tolerance of absent experts. We empirically find that different models may have different patterns of tolerance, which can be modeled with a neural network. After determining the memory allocation, we further calculate the update frequency based on the profiled I/O bandwidth - we always try to maximize the update frequency as long as the I/O operations does not block the computation.
We implement SwapMoE in two popular frameworks for large MoE models, including Huggingface Transformers (Wolf et al., 2019) and Tutel (Hwang et al., 2022). To evaluate the effectiveness of SwapMoE, we conduct experiments with Swin-MoE (Liu et al., 2022), Switch Transformer (Fedus et al., 2022) (SwitchT for short), and GPTSAN (gpt, 2023) models on both computer vision and natural language processing tasks, compared with the normal model inference scheme and strong baselines. For example, on text summarization tasks with Switch Transformer, SwapMoE can reduce the memory consumption from 14.2 GiB to 4.7 GiB (67% less), together with 50% latency reduction and a slight Rouge-2 score drop of 0.041.
These results have shown that SwapMoE enables the serving of large MoE models on resource-constrained edge devices with limited memory budgets.
We summarize our key technical contributions as follows:
  • We propose a novel MoE inference framework that supports efficient continuous serving of large MoE models under memory constraints. Our method enables deployment of off-the-shelf MoE models to edge devices with tunable memory.
  • We investigate several properties of MoE models, including the expert activation locality, per-sample expert importance, and layer-wise tolerance to absent experts, which can potentially benefit other works on MoE optimization.
  • We implement SwapMoE with popular inference frameworks and conduct extensive experiments with large MoE models on edge devices. The results have demonstrated the effectiveness of our approach on popular computer vision and neural language processing tasks.
2. Background and Motivation
2.1. MoE Preliminaries
Figure 1. Illustrations of MoE layers. MoE layers are usually used with Transformer architectures where the layer input is a sequence of tokens.
Increasing the capacity of a model can improve its performance (Kaplan et al., 2020). A prevailing method for scaling up models is using MoE, where certain model layers are replaced as parallel experts and are sparsely activated for inference based on their scores given by the gating network. In GPT-MoE (Artetxe et al., 2022), it is observed that some experts within the model exhibit a higher proficiency in English, while others demonstrate a greater aptitude for handling French inputs.
A typical MoE module is shown in Figure 1(b). Let $G(x)$ denote the output of the gating network for a given input $x$, and let $E_{i}(x)$ denote the output of the $i$-th expert network for the same input $x$. Then, the output $y$ of the MoE module can be expressed as:
(1) $y=\sum_{i=1}^{n}G(x)_{i}E_{i}(x),$
where $n$ is the number of expert networks in the MoE module, and $G(x)_{i}$ is the gating weight assigned to the $i$-th expert network for the input $x$. Most existing MoE models are based on the transformer structure (Vaswani et al., 2017). Therefore, the input $x$ in Equation 1 represents a token, and each layer of the MoE model receives multiple tokens as input. The gating weights $G(x)_{i}$ are typically calculated using a softmax function over a set of learned parameters. We can convert any layer of a deep learning model into an MoE module.
Among different MoE variants, the sparse MoE (Shazeer et al., 2017) is the most commonly used, in which only one or few (represented as $k$) experts are activated for each input. Specifically, in a sparse MoE with $k=1$, only the $i$-th expert with max $G(x)_{i}$ is activated for input $x$, and the output $y=G(x)_{i}E_{i}(x)$.
The MoE structure is often accompanied by the Transformer architecture (Liang et al., 2022; Zhang et al., 2022), in which the input of each MoE layer is a sequence of tokens and each token may choose different experts in one MoE layer, as shown in Figure 1(b).
2.2. Memory Constraints of Edge AI
In the past decade, the growth in memory capacity of mobile/edge devices has significantly lagged behind the advances in computing power. For example, from iPhone 4 in 2010 to iPhone 14 in 2022, the GPU and CPU performance of iPhone series has increased by nearly 100-1000 times (in terms of FLOPS) (John, 2016), but the RAM capacity has only increased by a factor of 8 (GadgetVersus, 2023).
Unlike cloud servers that can elastically expand the memory with distributed interconnected GPUs, the memory capacity of edge devices is usually a hard constraint. The AI models that exceed the memory constraints are usually infeasible to run on edge devices. For instance, the memory size of most edge devices today typically range from 1-8 GiB, while the model sizes of many state-of-the-art MoE-based language models and vision models are larger than 8 GiB.
Meanwhile, the memory constraints not only come from the hardware, but also from the operating systems and applications. Many edge devices have to serve multiple applications and each commercial application has multiple internal AI-based functions, so the memory budget for each AI model is further limited.
Figure 2. The memory footprint of each layer in SwitchT-32 (i.e. Switch Transformer with 32 experts per MoE layer). 89.42% of the memory is consumed by the MoE layers (even layers).
Therefore, reducing the memory footprint of MoE models is desirable to fit them into memory-constrained edge devices. As shown in Figure 2, most of the memory is consumed by the MoE layers, which is the bottleneck to optimize.
2.3. Problems of Conventional Solutions
Existing methods for reducing the resource overhead of MoE model inference include on-demand loading (Lane et al., 2016) (i.e. loading the parameters of MoE layers when they are needed and release afterwards) and MoE model pruning (Chen et al., 2022; Kudugunta et al., 2021) (i.e. cutting the less important experts permanently).
While the on demand loading method can reduce the GPU memory usage during MoE model inference without affecting model accuracy, it introduces significant latency overhead with each MoE layer parameter transmission via PCIe. In Figure 3, running MoE models with on-demand loading introduces 6.2x-8.9x higher latency. Meanwhile, the transfer rate of PCIe is approximately 10 GiB/s, whereas the large MoE model’s parameters are at least 1 GiB. Our measured IO bandwidth usage is 10.2 GiB/s, while the latency for one model inference is approximately 0.2 s. As a result, the transmission of a large amount of parameter data can significantly impede the computation process. As shown in Figure 4 and Figure 5, model parameters transmission takes up most inference time and significantly obstructs the model’s computational process and may deplete IO resources.
The pruning-based method directly reduces the model’s parameter and computational load, thereby lowering the GPU memory usage and inference latency during MoE model inference. However, it greatly compromises the model’s performance due to compromised the models’ capacity. For example, when utilizing expert pruning to reduce memory usage by 30% on SwitchT-32, the model’s accuracy decreased by 14%. SwapMoE combines the advantages of both methods - we try to reduce the latency overhead of memory saving while striving to maintain the model accuracy at the same time.
Figure 3. Naive on-demand expert loading reduces memory, but also results in huge inference overhead.
Figure 4. Latency breakdown of MoE model inference with layer-wise memory swapping. The transmission of model weights consumes the majority of the time.
Figure 5. Weight loading may block computation when running MoE with layer-wise memory swapping. Due to the large size of MoE layers and sparse computation, loading the weights of a layer is always slower than computing the layer, which slows down the inference even if the weights are loaded asynchronously.
2.4. Data Distribution Locality in MoE Serving
Data distribution locality is an important characteristic in many edge AI applications, which refers to the phenomenon that successive input samples are distributionally similar or correlated. The locality property is common in both language modeling and video analytic tasks that MoE models are mainly used for. Firstly, language modeling tasks are naturally subject to locality - each inference process of the model produces a token, and the successive tokens belong to the same sentence. Secondly, edge AI models are usually deployed in a fixed environment and used for serving an individual user or organization, the successive input samples (e.g. video frames, conversations) are semantically close to each other.
Since the experts in MoE models are designed to handle different data distributions, there exists an opportunity to cache the most relevant experts in the main memory at each time step, therefore reducing the memory consumption. The data distribution locality in edge applications further produces the chance to reuse the cached experts for successive input samples, which can reduce the overhead of parameter loading.
Such locality properties have been frequently and suceesfully exploited for edge AI inference acceleration (Wakatsuki et al., 2021; Kong et al., 2023). For example, one can skip or simplify the inference process of some video frames based on previous inference results in video analytic tasks. These approaches inspired our design, but are orthogonal to our work. In this paper, we aim to exploit the locality to optimize memory-constrained serving of MoE models, including large language models. Both the memory aspect and MoE models are rarely discussed in prior locality-based optimizations.
3. Our Design: SwapMoE
Figure 6. The workflow of SwapMoE. Given an off-the-shelf MoE model and a memory-constrained edge device, we satisfy the constraint by executing the model with a smaller set of experts (Virtual Experts). The Virtual Experts are selected, used and updated seamlessly at runtime, and the memory allocation for the experts is determined at offline.
We introduce SwapMoE to achieve efficient and accurate MoE inference under tunable memory budgets. The main idea is to perform MoE inference with a subset of “Virtual Experts” in memory (which reduces the memory footprint) and seamlessly update the Virtual Experts according to their relevance to current data distribution (which maintains accuracy and latency). The key technical challenges to be addressed in SwapMoE include:
  1. (1)
    How to select and use the Virtual Experts in an off-the-shelf MoE model, so that the MoE model can produce accurate prediction with a small set of Virtual Experts.
  2. (2)
    How to dynamically manage the Virtual Experts via seamless memory operations that can minimize latency/memory overhead at runtime.
3.1. Overview
SwapMoE addresses the challenges with various techniques in a two-phase holistic design, as shown in Figure 6.
In the online phase, the job of SwapMoE is to efficiently identify, update, and use a subset of experts (Virtual Experts) for memory-constrained MoE inference. We utilize a lightweight magnitude-based method to estimate the importance score of each expert, and maintain the most important experts in memory (i.e. Virtual Experts). These experts are used to handle the whole inference process, which will be discussed in Section 3.2. Furthermore, SwapMoE needs to dynamically updates the Virtual Experts to maintain important experts seamlessly (without impacting the latency), which will be discussed in Section 3.3.
The job of the offline phase is to obtain an optimal memory plan for the online phase to facilitate efficient and accurate MoE inference. To achieve this, we first obtain several important performance metrics and models through fine-grained expert profiling, including memory footprint, compute latency, loading latency, and model accuracy, which will be discussed in Section 3.4. The profiling needs to be conducted only once for a specific device. Then the profiling-guided expert planner utilizes the established relations to derive the optimal memory plan (layer space allocation and expert I/O frequency) for Virtual Experts based on a given memory budget, which will be discussed in Section 3.5.
3.2. Importantance-aware Virtual Experts Selection & Inference
To maintain accuracy and latency with Virtual Experts-based MoE inference, we must wisely select the most important experts and update them according to the data distribution to maintain accuracy with Virtual Experts-based MoE inference. To ensure that the Virtual Experts remain up-to-date and consist of the most important experts for the current data distribution, the first question is how to determine the importance of each expert at runtime. To address this problem, we introduce the concept of the expert importance score.
The expert importance score is an input-dependent metric, denoted as $importance(E_{i},X)$, that approximates the positive contribution of the expert $E_{i}$ to the inference process of the input $X$. Keeping the experts with higher importance scores in the Virtual Experts are expected to produce more precise predictions with the MoE model.
[…]
Based on the mechanism of MoE in Equation 1, we notice that each expert $E_{i}$ influences the output of the MoE layer by adding its output $G(x)_{i}E_{i}(x)$ to the output. Naturally, we speculate that the key factor determining the importance of $E_{i}$ is the value of $||G(x)_{i}E_{i}(x)||$, where |||||$\cdot$||| | ⋅ | | represents the magnitude. However, we cannot directly calculate $||G(x)_{i}E_{i}(x)||$ for each expert during the inference process, because the non-Virtual Experts experts are not available in the main memory to do the inference of $E_{i}(x)$. As a result, we need to approximate the value of $||G(x)_{i}E_{i}(x)||$ without calculating $E_{i}(x)$.
After analyzing Equation 1, we approximate $||G(x)_{i}E_{i}(x)||$ using $||x||*||G(x)_{i}||*||E_{i}||$ to determine the importance of experts, where $||E_{i}||$ is the parameter magnitude of $E_{i}$ that can be pre-computed offline. The rationale behind this approximation is that the larger $||x||$, $||G(x)_{i}||$ and $||E_{i}||$ typically lead to higher $||G(x)_{i}E_{i}(x)||$ in high-dimention matrix multiplication, and vice versa. Therefore, the final formula to calculate the importance of experts is
(2) $importance(E_{i},X)=\sum_{x}^{\mathcal{X}_{i}}||x||*||G(x)_{i}||*||E_{i}||,$
where $\mathcal{X}_{i}$ is the set of tokens passed to expert $E_{i}$ during the inference process of $X$. This score only involves simple magnitude computation thus can be calculated efficiently.
Next, we explain how SwapMoE perform inference only with the selected Virtual Experts in main memory. Once Virtual Experts has been selected, we can calculate the output of the MoE layer, as defined by Equation 1.
Figure 7. (a) Original gating of MoE: all experts may by used for inference; (b) Ours Masked Gating: only Virtual Experts will be used.
As shown in Figure 7, we use Masked Gating to redirect all inference requests to Virtual Experts. Specifically, once the computation of $G(x)$ in Equation 1 is completed, we achieve the flow of all tokens’ inference requests to Virtual Experts by assigning scores corresponding to unimportant experts as $-inf$:
$G(x)_{i}=-inf,\text{if i not in Virtual Experts},$
then the function $G(x)$ will be adjusted through the softmax operation. In essence, by performing the max operation on $softmax(G(x))$, none of the unimportant experts will be selected.
3.3. Seamless Virtual Experts Update
To run Virtual Experts-based MoE inference efficiently, we need to seamlessly perform the expert updates without latency overhead. Once we have the importance scores for all experts, we can update Virtual Experts accordingly at runtime by loading the important experts into the main memory and the unimportant experts out. In our design, the number of experts in the Virtual Experts is fixed, so we only need to load the parameters of new experts to overwrite the old experts.
The most critical question that arises in expert loading is the additional I/O cost, which may cause idle waiting time before computation and significantly slow down the inference process. We introduce two techniques to reduce this overhead, including amortized expert loading and asynchronous expert loading, as shown in Figure 8 (ii).
Amortized expert loading.
Figure 8. Different loading strategy of Virtual Experts. i) synchronously updating all experts after single sample inference; ii) amortized asynchronous expert loading across different samples.
According to Equation 2, the importance scores of all experts in all MoE layers may be updated for each input sample. One nature way is loading experts for all layers and every sample during inference. This method results in optimal accuracy, but the overall system latency is unacceptably high due to the large number of experts in the MoE model. For instance, in the SwitchT-16 model (Switch Transformer with 16 experts per MoE layer), if we update 50% of the experts, approximately 5GB of data needs to be transferred (typical data transfer speed of GPUs is approximately 10 GiB/s), resulting in high latency. Therefore, we adopt a fixed frequency for updating, which is determined by the update frequency set in the profiling phase.
However, simply reducing the frequency of updates is not enough. If we do the loading of expert periodically, the samples that undergo expert loading would suffer from high inference latency due to the extensive data transmission and limited bandwidth between the main memory and external memory (or disk). While other samples that do not undergo expert loading will have lower inference latency. Thus, we further distribute this periodic overhead of expert parameter transmission across different input samples to fully utilize the I/O bandwidth. Specifically, in each update cycle that spans $k$ samples, we evenly partition the pending experts (that are determined to load into Virtual Experts) to $k$ samples. Within every $\frac{k}{\text{\# of moe layers}}$ we just load experts of one MoE layer. In this way, we can distribute the cost of expert loading to samples.
Asynchronous expert loading. The amortized loading strategy can reduce the size of per-sample data transmission, but it will still increase the inference latency due to the additional I/O waiting time. Fortunately, we observe that replacing experts in Virtual Experts does not affect the inference computation process. Therefore, we adopt a computing while transferring strategy, where the actual replacement of the new experts is performed asynchronously during the inference computation of old experts. Once the loading process of new experts is done, the inference process is notified to use the new experts and expire the old experts. While the I/O cost of expert loading is not changed in this asynchronous design, the delay in the overall inference latency is mitigated by the concurrent nature of transferring and computing, ensuring the efficient execution of the model.
3.4. Fine-grained Expert Profiling
To facilitate efficient and accurate model inference, it is necessary to know the performance of Virtual Experts given a specific hardware and configuration and determine what kind of configuration will lead to better perfromance. We conduct fine-grained expert profiling in advance, gather information related to hardware memory usage, inference latency, accuracy, and IO bandwidth, and establish the relationship between Virtual Experts configurations and performance, Specifically, configurations are:
[…]
(3) num_experts ={#$=\{\#experts_{1},\#experts_{2},...,\#experts_{L}\},$
where $frequency$ represents the update frequency of the experts in the core, indicated as the number of inputs between each pair of updates. #$\#experts_{i}$ denotes the number of experts to be retained in the core for the $i$-th layer.
Problem Formulation. The primary objective of offline planning is to identify the optimal configuration $\hat{\text{config}}$ that meets memory constraints while maximizing the model’s accuracy and minimizing inference latency. Formally, the process can be described as follows:
$maximize$ ${E_{\text{accuracy}}(\text{config})},$
$minimize$ ${E_{\text{latency}}(\text{config})},$
(4) s.t. $E_{\text{memory}}(\text{config})\leq{LIMIT}_{\text{memory}},$
where $E_{\text{accuracy}}$, $E_{\text{memory}}$, and $E_{\text{latency}}$ are used to estimate the accuracy, memory footprint, and inference latency of a model under a specific configuration. ${LIMIT}_{\text{memory}}$ represent the constraints on main memory.
3.4.1. Memory Footprint and Latency Estimation
In order to estimate the MoE model’s inference latency and memory, we need to conduct a more detailed profiling of the experts, encompassing the inference memory footprint, latency, and loading time. This is crucial as the overall cost is composed of numerous experts within the Virtual Experts.
Expert memory footprint: We conducted a detailed profiling of the memory footprint for the inference of each individual expert within the MoE layer. This includes the memory occupied by the parameters of each expert and the memory occupied by the activations generated during inference computation. Expert inference latency: The expert inference latency encompasses both data transmission and computation. We profiled the data transmission and computation time for each expert. Given that the parameters of each expert within a layer are identical, we only need to profile one expert per layer. Expert loading time: The loading time for each expert refers to the time taken for expert parameter transmission. As expert loading involves the transmission of data in I/O, we also need to profile the I/O resources to ensure that computation does not get blocked during expert loading.
With expert-level profiling data, we can obtain the whole model performance, where the computational load during model inference is mainly related to the number of experts in the configuration: the more experts there are, the larger the inference latency and memory footprint will be, and vice versa. While the expert’s inference latency and memory differ across different hardware, we only need to model it once for a given hardware to obtain $E_{\text{memory}}$, and $E_{\text{latency}}$.
3.4.2. Accuracy Influence Modeling
Directly measuring the accuracy of a configuration on the target device is time-consuming because it requires running the MoE model under the configuration multiple times. When the configuration space is large, measurement is impractical. Predicting the accuracy is a promising option, but the complex correlation between the configuration and the accuracy makes it challenging, which have been studied in previous work (Zhang et al., 2021). While it is more convenient to predict accuracy indirectly through modeling, the accuracy of the model itself is inherently unexplainable.
To obtain $E_{\text{accuracy}}$, we first need to collect a small amount of profiling samples from the target device that can reflect the data distribution of the deployment scenario. Profiling samples must be labeled to measure the accuracy of SwapMoE under different configurations. Since the deployed models will be used in the target environment for a long time, it is feasible to collect such profiling data. The data labeling can be done manually or with an oracle model.
Next, we generate a set of random configurations of Virtual Experts. For each configuration, we use SwapMoE under the configuration to perform model inference with the profiling dataset. We collect the corresponding $E_{\text{accuracy}}$. Note that when collecting $E_{\text{accuracy}}$, SwapMoE is operated by the expert scheduler, which can refer to Section 3.2 and Section 3.3.
Finally, we use deep learning to fit the relation between Virtual Experts configurations and the model accuracy. We use a small DNN (two-layer linear network work) to model the relation because it is lightweight and sufficient. The model minimizes the residual sum of squares between the actual model $E_{\text{accuracy}}$ by the lightweight DNN. The training of this DNN is exceptionally fast, with very low computational requirements, typically completing in just a few minutes (on a Jetson Nano).
3.5. Profiling-guided Memory Planning
Section 3.2 and Section 3.3 have shown how the Virtual Experts are selected and used for inference at runtime. However, many questions remain unanswered. For example, how to distribute Virtual Experts across different layers, how to allocate limited memory size to different layers, and how to make the best use of limited memory bandwidth and enabling frequent updates of experts without blocking computations, etc. These questions are crucial to satisfy the memory constraint and minimize accuracy loss. They are addressed through a profiling-guided memory planning, which using profiling information obtained from Section 3.4.
3.5.1. Expert I/O Frequency
The update frequency affects the usage of IO bandwidth, and if too much data need to be transferred through IO, it will cost a long time and block the computation. In other words, when the experts in Virtual Experts are no longer important for the current sample, we need to promptly replace them with the latest important experts. A higher update frequency is naturally preferable, but increased frequency can escalate hardware IO resource consumption, potentially even obstructing inference computation and leading to increased latency. At the same time, because a higher frequency is more beneficial for dynamically maintaining the most important experts, it is important to increase the frequency as much as possible without blocking the computation. Our strategy involves testing from a low update frequency to a high update frequency until we identify the inflection point at which the update frequency affects the inference latency, allowing us to select the optimal update frequency.
3.5.2. Layer Space Allocation
To allocate the limited memory budget to different layers, allow layers with more memory to utilize more Virtual Experts, find the optimal configuration ($\hat{\text{config}}$) that maximizes the model accuracy while satisfying given memory constraints, we utilize expert planner to obtain the optimal memory allocation scheme based on the previously obtained performance model obtained from $E_{\text{accuracy}}$, $E_{\text{memory}}$, and $E_{\text{latency}}$ in Section 3.4.
A naive method to find the optimal configuration is to iterate over all possible configurations and keep track of the best-performing one. However, we found that this approach does not fully leverage the modeled functions to find the optimal configuration due to the enormous search space. For instance, the search space in a 12-layer SwitchT-16 would be $12^{16}$.
Consequently, we decided to employ a genetic algorithm for the search process. Specifically, we initialize a set of configurations randomly, and iteratively update them based on their performance metrics estimated with $E_{\text{accuracy}}$, $E_{\text{memory}}$ and $E_{\text{latency}}$. In each iteration, we randomly change one parameter in each configuration and create new configurations by exchanging or averaging existing configurations. The configurations that violate resource constraints or yield suboptimal performance are removed. We observed that certain layers in the model may have a more significant impact on model performance. The genetic algorithm can perceive this characteristic and preserve experts that have a greater influence on performance. In Section 5, we will illustrate some configurations found by the algorithm.
4. Implementation
We implemented SwapMoE atop two popular MoE frameworks, Tutel (Hwang et al., 2022) for Swin-MoE, and Huggingface Transformers (Wolf et al., 2019) for Switch Transformer (Fedus et al., 2022) and GPTSAN (gpt, 2023), with 4k LoC added. Tutel uses customized operators for MoE layers, which employ token padding techniques to enhance load balancing and shape alignment between experts. Since the main focus of this paper is not to address the impact of padding, we directly adopted the token padding techniques used in Tutel.
Efficient implementation of Virtual Experts. To minimize the additional memory footprint caused by experts’ swapping in and out of main memory, we abstract Virtual Experts as a pointer concept. Virtual Experts remains stored in main memory, a portion of memory that is pre-allocated and does not require reallocation. The expert weights contained in Virtual Experts are those to which it points. When loading experts to Virtual Experts are necessary, we utilize the torch.copy_ API in PyTorch for transferring expert weights, thereby avoiding the need for additional main memory allocation. This process effectively updates the experts to which Virtual Experts points.
Genetic algorithm configurations. We provide the parameters used in the genetic algorithm to search the optimal Virtual Experts configuration. The population size is 50, the mutation rate is 0.5, the crossover rate is 0.01, and the number of generations is 5000. To ensure feasible results, we set the lower bound of the search space to 1 and the upper bound to the total number of experts in the MoE layer.
Profiling data processing. During the implementation of SwapMoE, we utilized 10% of the test data as profiling data. To have a sufficiently long profiling data set, we replicated this 10% of data multiple times and concatenated the beginning with the end. It is worth mentioning that label is not required for all the data in the online phase.
5. Evaluation
Figure 9. End-to-end latency achieved by SwapMoE and baselines under different memory budgets on Jetson AGX. SwapMoE achieves similar latency with the pruned compact models while with much more parameters in use.
(a) Jetson AGX
(b) Jetson Nano
Figure 10. The accuracy achieved by SwapMoE and the pruning baseline under different memory budgets on (a) Jetson AGX and (b) Jetson Nano. SwapMoE achieves significantly higher accuracy than the baseline in almost all cases.
We evaluate SwapMoE to answer the following questions:
  1. (1)
    Is SwapMoE able to provide better latency-accuracy and memory-accuracy tradeoffs than baselines?
  2. (2)
    Is SwapMoE able to find the optimal configuration that meets the constraints while maximizing accuracy?
  3. (3)
    What about the robustness of SwapMoE against different numbers of experts, input data locality, and hardware?
  4. (4)
    How much does each component contribute to the effectiveness of SwapMoE?
  5. (5)
    What about the overhead of SwapMoE?
5.1. Experimental Setup
Edge environments. We use two edge devices: a Jetson Nano and a Jetson AGX ORIN. The batch sizes are all set to 1 as in common edge-side continuous serving scenarios. We use different latency and memory budgets to simulate different resource constraints.
Tasks, datasets and models. We evaluate the performance of SwapMoE on three common DL tasks:
Object detection aims to detect objects in an image, predicting the object bounding boxes and categories. We chose Cascade with Swin-MoE (Liu et al., 2022) base backbone as the detection model, and trained 32 experts Swin-MoE models with the Shift dataset (Sun et al., 2022) for evaluation. The performance of the detection model is measured by mean average precision over Intersection over Union (mAP) with IoU threshold at 0.5, where higher mAP means better.
Summarization aims to summarize long texts into short texts. We select the most popular summarization model, Switch Transformer (Fedus et al., 2022) (16, 32, and 64 experts per layer, denoted as SwitchT-16, SwitchT-32 and SwitchT-64) and use the samsum dataset (Russakovsky et al., 2015) and report Rogue-2 accuracy, where higher Rogue-2 accuracy means better.
Language modeling aims to predict the next word given the previous words. We use GPTSAN (Artetxe et al., 2022) as the base model and use the Wikipedia-japanese (wik, 2023) dataset for evaluation. The performance of the language model is measured by perplexity, where lower perplexity means better.
The pre-trained weights of the Switch Transformers and GPTSAN are obtained from the official Huggingface Transformer repository. The Swin-MoE model were trained by ourselves according to the source code and the default configurations of their papers. All datasets can be downloaded from public websites.
Baselines. The basic baseline is original MoE model. To show the superiority of SwapMoE, we also compare it with the following baselines: ‘Pruning’ (keeping a certain portion of experts with the largest $||E_{i}||$ in each layer, no switching in/out), ‘On demand’ (Huang et al., 2023) (keep a certain portion of experts in main memory, load the requested expert from memory on demand).
5.2. Overall Runtime Performance
Method Constraints Achieved performance
Memory budget (GiB) Memory (GiB) Latency (s) mAP (%)
Original MoE \ 1.99 0.25 52.60
vExperts 0.9 0.94 0.13 51.23
1.2 1.11 0.15 51.65
1.5 1.44 0.19 52.26
1.8 1.79 0.23 52.58
(a) Swin-MoE
Method Constraints Achieved performance
Memory budget (GiB) Memory (GiB) Latency (s) Rogue-2
Original MoE \ 4.08 0.82 0.24
vExperts 2.0 1.64 0.45 0.21
3.0 2.74 0.65 0.22
3.5 3.58 0.76 0.23
4.0 3.99 0.80 0.23
(b) SwitchT-16
Table 1. Performance of SwapMoE with different MoE models under different memory constraints.
Our approach offers good memory-latency tradeoffs. As shown in Figure 9, in terms of latency, our approach and the ‘Pruning’ method exhibit very similar memory-latency trade-offs. This indicates that our approach, while ensuring minimal model accuracy degradation, enables the model to occupy less memory and achieve reduced inference latency. Although the ‘On demand’ method can significantly reduce memory usage, its cost comes in the form of high latency overhead.
Our approach also offers good memory-accuracy tradeoffs. As shown in Figure 9(a) and Figure 9(b), SwapMoE creates a trade-off space that accommodates various resource usages. Higher resource utilization leads to better model performance and vice versa; and, even if resource consumption reduces, SwapMoE still maintains the model’s performance. For example, in summarization task, SwapMoE reduces memory usage and latency by 37% and 18% with only 0.012 Rouge-2 degradation on Jetson AGX.
It is worth noting that SwitchT-64 cannot be directly deployed on Jetson Nano, as its hardware has a maximum memory support of 4GB, while the inference demand of SwitchT-16 exceeds 4GB. Nevertheless, in Figure 9(b), our approach significantly outperforms the ’Pruning’ baseline. For instance, on Jetson Nano, when reducing the memory footprint of SwitchT-64 from 14GiB to 2.6GB, SwapMoE achieves an accuracy that is 175% higher than the baseline.
Furthermore, SwapMoE is a general method applicable to various tasks, including object detection, summarization, and language modeling tasks. In detection tasks, significant reductions in memory and latency resulted in minimal accuracy loss. However, in language modeling, reducing memory and latency introduced some deterioration in perplexity. This is because in computer vision (CV) tasks, the pixel values between consecutive input images are also continuous, thus exhibiting stronger locality. In contrast, for natural language processing (NLP) tasks, consecutive inputs are only semantically similar, such as in tone or knowledge domain, and therefore do not exhibit the same level of locality as in CV tasks.
SwapMoE also outperforms the baselines, demonstrating its effectiveness. Under the same memory budget, it achieves better performance than the baselines. For example, in SwitchT-32, when the memory usage is 4.7GB, the Rugue score of SwapMoE is 0.232, which is 8.76% higher than the ’Pruning’ baseline. This is because we select the most important experts based on the characteristics of input samples, rather than simply pruning the MoE model into a smaller but static model. However, we cannot surpass the on-demand approach because it does not alter the model’s output. However, its high latency makes it unsuitable for edge scenarios where both latency and memory are constrained.
5.3. Offline Planning Performance
SwapMoE can find optimal configurations that satisfy given constraints while maximizing model performance, as shown in Table 1. For example, in the object detection task, when given a memory budget of 1.5 GiB, the configuration found by SwapMoE allows the model to achieve an actual 1.44 GiB peak GPU memory usage and 0.19 s inference latency, satisfying the constraints while only experiencing a 0.32% decrease in accuracy compared to the original MoE model.
Furthermore, it is observed that as the constraints become looser, SwapMoE can achieve higher accuracy, and conversely, tighter constraints result in lower accuracy. It found optimal configurations under various constraints, respectively, to maximize the utilization of resources. For example, in the summarization task, when the memory budget changed from 3.5 GiB to 4 GiB, the actual peak GPU memory and inference latency of the model during runtime changed from 3.58 GiB and 0.76 s to 3.99 GiB and 0.80 s, respectively. The accuracy also changed from 0.23 to 0.24. However, our method is not always able to satisfy the budget. For example, in the summarization task with SwitchT-16, when the memory budget is set to 3.5 GiB, the actual peak memory usage and inference latency achieved by the model are 3.58 GiB and 0.76 s, respectively. This situation arises because the inference performance of the model itself is difficult to predict accurately (Zhang et al., 2021), leading to inaccurate performance modeling.
(a) SwitchT-16
(b) Swin-MoE
Figure 11. Distribution of allocated memory across layers under different memory budgets in SwapMoE.
Additionally, we find that the distribution of Virtual Experts across different layers differs in different tasks. In the language modeling task, SwapMoE tends to maintain more experts in the middle layers, as shown in Figure 11 (a). This suggests that for language models, the intermediate MoE layer has a greater impact on model inference, as the intermediate layers have a more significant influence on logical processing. In contrast, in the object detection task, SwapMoE maintains more experts in the early layers, as shown in Figure 11 (b). This indicates that for visual models, the foremost MoE layer has a greater impact on model inference, as the preceding layers have a more substantial influence on image feature extraction. For example, when the constraints are looser in the object detection task, the total number of SwapMoE increases, and the number of SwapMoE in the early layers increases. This indicates that the early layers are more important for the object detection task.
5.4. Robustness Analysis
Figure 12. Robustness of SwapMoE under different usage scenarios.
Refer to caption (a) Different data locality
(b) Different hardware
Refer to caption Refer to caption
Figure 12. Robustness of SwapMoE under different usage scenarios.
Figure 13. Ablation study: performance of SwapMoE with different components replaced.
In this experiment, we analyze the robustness of SwapMoE across different usage scenarios.
Different number of Experts. SwapMoE can be used for MoE models with different numbers of experts. As shown in Figure 9(a), SwapMoE finds good performance-resource tradeoffs between Switch Transformers (16, 32, and 64) with different resource usage. And SwapMoE can reduce the resource consumption of SwitchT-64 to a level similar to that of SwitchT-32, while maintaining comparable accuracy. This demonstrates that our method can effectively prune large MoE models into well-performing smaller models based on resource constraints, eliminating the need to store multiple model sizes.
Different data locality. SwapMoE can be used with input data with different patterns. As shown in Figure 11(a), when SwapMoE works with lower temporal locality data (faster video with 3x video speed and random shuffled video frames), it is still able to find good performance-resource tradeoffs. And as the temporal locality of the data decreases, the performance of SwapMoE deteriorates. We observed that SwapMoE performs best on normal videos, followed by faster videos with 3x video speed, and finally random shuffle video frames. This is because as the temporal locality of the data decreases, the importance of experts in the MoE model becomes more unstable, resulting in poorer performance of SwapMoE under the same resource usage conditions. In other words, to achieve the same model performance, SwapMoE requires more resource usage on data with lower temporal locality.
Different hardware. SwapMoE can be used on devices with different I/O bandwidths. As shown in Figure 11(b), SwapMoE can find good performance-resource tradeoffs on the Jetson AGX device. Additionally, we implement two versions of SwapMoE on Jetson AGX. (1) CPU-GPU: swapping experts between CPU memory and GPU memory. (2) GPU-Storage: swapping experts between GPU memory and external storage. The experimental results demonstrate that despite lower storage bandwidth, SwapMoE is still able to find excellent resource-performance trade-offs. This is because SwapMoE only requires to swap a small number of experts for each sample at runtime. A small bandwidth between the main memory and the external memory is sufficient.
5.5. Ablation study
In Figure 13, we show the performance of SwapMoE when replacing one component from the system. The experimental results indicate that the absence of any component in SwapMoE leads to performance degradation. In other words, all components contribute to the performance of SwapMoE.
(1) ‘Simple scheduling’: change the scheduling in SwapMoE component, remove amortized updating component, and calculate expert importance score where experts with more tokens to inference have higher scores. The perplexity of SwapMoE is better than ‘simple scheduling’ because the importance score obtained by expert scheduler is more accurate via approximating the outputs of experts, instead of simplely counting the number of tokens dispatched to each expert. Therefore, SwapMoE choose the important experts that are more suitable for the current input samples. Furthermore, ”simple scheduling” currently lacks amortized updating, which results in higher latency. This illustrates the necessity of amortized updating, as it can significantly reduce inference latency through the process of amortizing computations.
(2) ‘Simple planning’: change the planning in SwapMoE component to ‘simple planning’ and adopt an even Virtual Experts distribution, meaning that each MoE layer has an equal number of Virtual Experts, and then selects the best from these search spaces, where search space is very limited and does not encompass the optimal solution. The perplexity of SwapMoE is superior to ‘simple planning’ because the expert planner utilizes genetic search to find the optimal configuration for SwapMoE from a large configuration space. In contrast, ‘simple planning’ cannot find a worse configuration than SwapMoE due to the suboptimal search process.
Memory External Memory IO Overhead (MiB/s)
Constraint (GiB) Consumed (GiB) Peak Mean
1.2 1.02 40 20
1.5 1.02 36 27
1.8 1.02 13 11
Table 2. The runtime overhead of SwapMoE.
5.6. Overhead
We report the overhead of SwapMoE in the object detection task with the Swin-MoE model.
Offline planning overhead. The offline planning phase includes performance modeling and optimal configuration generation. Performance modeling only needs to be done once for a MoE model, which takes about 20 minutes. Optimal configuration generation needs to be done with each different resource constraint, which takes about 5 seconds.
Runtime overhead. The runtime overhead includes calculating expert importance score, expert swapping, and different expert request handling strategies. As shown in Table 2, the peak IO overhead of SwapMoE is about 20 MiB/s, which is negligible compared to the IO bandwidth between the main memory and the external memory (e.g., 10-30 GiB/s for GPU-CPU over PCIe and 300-600 MiB/s for CPU-SSD). This is because the design of SwapMoE enables us to swap only a small number of experts. The “External Memory Consumed” means the space needed in the external memory (CPU memory or storage) to store the weights of experts. We store the original MoE model parameters in external memory to reduce the usage in main memory.
6. Related Work
Systems optimization for MoE model serving. Common techniques for optimizing MoE model serving include offloading and swapping memory. Huang et al. (Huang et al., 2023) propose to swap the experts from GPU memory to CPU memory to reduce the memory consumption of MoE models, incurring high latency overhead. SE-MoE (Shen et al., 2022) utilizes Ring Memory offloading to reduce GPU memory consumption of MoE models. However, these methods are not suitable for resource-constrained devices with dynamic latency and memory constraints. The proposed SwapMoE is the first work to optimize MoE model serving on resource-constrained devices with dynamic memory constraints.
Optimization for dynamic DL model serving. MoE is a type of dynamic neural networks. Many approaches are introduced to enable or enhance such dynamic DL model serving in general. Nimble (Shen et al., 2021) is a system that optimizes, compiles, and executes dynamic neural networks on multiple platforms by using a dynamic type system and a lightweight virtual machine runtime. Model scaling approaches (Fang et al., 2018; Han et al., 2021; Wen et al., 2023) propose to adjust the model size/architecture on edge devices to meet different resource constraints. Remix (Jiang et al., 2021) proposes to use multiple models and dynamically switch between them during the inference process. Besides serving dynamic models, researchers have also attempted to slice the static model to dynamic components to achieve different resource-performance tradeoffs (Hou et al., 2022; Zhang et al., 2020). As compared with these approaches, our design is fundamentally different because it is based on the unique structure and characteristics of MoE.
Efficient design of MoE models. Many existing approaches study the efficiency problem of MoE models from the model design perspective. GShard (Lepikhin et al., 2020) scales up Transformer with MoE and improves the quality and efficiency of multilingual machine translation. Task-MoE (Kudugunta et al., 2021) extract subnet from a large MoE model by a task-level routing strategy. Chen et al. (Chen et al., 2022) propose to prune non-professional experts according to the downstream task for efficient MoE deployment. MPoE (Gao et al., 2022) proposes to build a parameter-efficient MoE architecture by enforcing parameter sharing between the experts. AutoMoE (Jawahar et al., 2022) utilizes neural architecture search to automatically design more efficient MoE models. These methods need to modify and retrain the MoE model to reduce resource consumption. They are orthogonal to our approach and can use our method to further reduce resource consumption.
7. Conclusion
This paper addresses the challenges of deploying MoE models to resource-constrained edge devices. We propose a framework called SwapMoE that adaptively reduces the inference costs of MoE models based on the memory constraints while preserving the model accuracy. Experimental results have demonstrated that our method can significantly reduce the inference costs of MoE models with reasonable accuracy degradation. Our approach creates a nice space of resource-accuracy trade-off of SoTA large MoE models.